OpenStudy (anonymous):

When planning his crops, Farmer Joe knows that his 5 acres can support corn and strawberries. He wants to make $5,500 from his crops. He can make $1,000 per acre of corn (the variable c) and $1,500 per acre of strawberries (the variable s). Which equation below would be a constraint in his system of equations? c + s = 5 c − s = 5 1,500c + 1,000s = 5,500 1,500s − 1,000c = 5,500

OpenStudy (danjs):

hi

OpenStudy (danjs):

c = acres of corn s = acres of strawberries . .

OpenStudy (danjs):

Both the amount of money he makes, and the total Crop Acres are related to ..C and S

OpenStudy (anonymous):

okay I'm following

OpenStudy (danjs):

Total Acres = Acres Corn + Acres Strawberries Total $ = (Acres Corn)*($corn per acre) + (Acres Strawberries)*($Straw per acre)

OpenStudy (danjs):

The problem gives, the total $ he wants, the price per acre of each crop, and the total acres of the farm

OpenStudy (anonymous):

yeah I'm following

OpenStudy (danjs):

putting those values in gives... 5 = C + S $5500 = $1000*C + $1500*S

OpenStudy (danjs):

5 acres equals acres of corn plus acres straw.. 5500 equals $1000 times acres of corn plus $1500 times acres of straw

OpenStudy (danjs):

that is your system of equations

OpenStudy (anonymous):

so the answer would be c+s=5 ?

OpenStudy (danjs):

right

OpenStudy (anonymous):

thank you so much